OpenStudy (anonymous):

Solve P = 2w + 2h for w. Please help:)

hartnn (hartnn):

subtract 2h from both sides.

OpenStudy (shaik0124):

p-2h/2=w

OpenStudy (anonymous):

P-2h=2w

OpenStudy (shaik0124):

now divide by 2 both sides

hartnn (hartnn):

now divide 2 both sides.

OpenStudy (anonymous):

so 2h divided by 2 and 2w divided by 2?

hartnn (hartnn):

the whole P-2h will be divided by 2, isn't it ?

OpenStudy (anonymous):

Ohhhh yeahh i forgot!

OpenStudy (anonymous):

so is it w= p-2h divided by 2? as the final answer ?

hartnn (hartnn):

yup :)

OpenStudy (anonymous):

Yay:) thanks again!

hartnn (hartnn):

welcome again :)